# DReSS: Data-driven Regularized Structured Streamlining for Large Language Models

### v2

@@ 4.6 Minimal Overhead
We evaluated the perplexity of each method on LLaMA2-7B using varying amounts of data, keeping all other conditions consistent with Section 4.2 and only change the data size from 500 to 8,000. As shown in Table 3, when only 1,000 samples were used for regularization and RFT, DReSS outperformed the other methods that used 4,000 samples. This further highlights the effectiveness and efficiency of applying regularization prior to pruning, as it transfers critical information from the pruned segments in advance, enabling DReSS to deliver strong performance with minimal data overhead.
Table 3: The impact of the amount of data on model performance.
Data Size 500 1000 2000 4000 8000 LLM Surgeon 8.14 7.38 6.55 5.89 5.56 SliceGPT 8.63 7.55 6.82 6.07 5.74 SLEB 13.26 10.24 9.67 8.79 7.83 DReSS 6.75 5.86 5.61 5.48 5.43
@@ 4.7 Ablation Study
Effectiveness of Regularization: As shown in Table 4, both random pruning and full parameter regularization significantly increase perplexity on WikiText-2 and reduce average accuracy at a 25% pruning ratio. This highlights the importance of applying regularization exclusively to the pruned components. Additionally, RFT after pruning has minimal impact on model performance, indicating that the RFT process is optional.
Table 4: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘FPR’ refers to full parameter regularization, ‘R’ refers regularization. ‘w/o’ denotes the removal of specific components.
Model Method Setting PPL($\downarrow$) $\Delta$ AVG_ACC $\Delta$($\downarrow$) LLaMA2-7B DReSS 5.86 0.00 61.54 0.00 w FPR 12.68 +6.82 50.26 -11.28 w/o R 22.38 +16.52 47.25 -14.29 w/o RFT 5.97 + 0.11 59.56 -1.98 LLaMA2-13B DReSS 5.12 0.00 68.84 0.00 w FPR 10.53 +5.41 53.28 -15.56 w/o R 18.94 +13.82 50.17 -18.67 w/o RFT 5.39 + 0.27 67.96 -0.88
Table 5: The perplexity on WikiText-2 under different regularization and RFT data ratio using LLaMA2-7B at a pruning ratio of 25%.
Ratio (R/F) 4:1 3:1 2:1 1:1 1:2 1:3 PPL($\downarrow$) 5.89 5.86 6.07 6.73 7.45 8.62
Table 6: The perplexity of different $\lambda$ on LLaMA2-7B.
$\lambda$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $5\times 10^{-3}$ $10^{-2}$ $5\times 10^{-2}$ $10^{-1}$ PPL($\downarrow$) 7.47 6.04 5.86 5.93 6.12 7.26 10.55
Data Ratio of Regualrization and RFT: With all other settings consistent with Section 4.2, we only varied the data ratio, and the results are shown in Table 5. It can be observed that when the ratio is 3:1, the model achieves the lowest perplexity. When the ratio is 4:1, the increase in perplexity is less pronounced, but when the ratio is 1:1, the perplexity increases significantly by 15%. These results underscore the importance of the regularization process relative to the RFT applied after pruning.
Trade-Off Between Language Modeling and Regularization: Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$. As shown in Table 6, when $\lambda=10^{-5}$, the regularization loss is too small to effectively shift important values, leading to a significant increase in perplexity after pruning. In contrast, when $\lambda=10^{-1}$, excessive regularization compromises the model’s language modeling ability, causing a notable performance decline post-pruning. Optimal performance is achieved at $\lambda=10^{-3}$, demonstrating the importance of balancing language modeling loss and regularization loss for effective pruning.
@@ 5 Conclusion
In this paper, we introduce a novel pruning paradigm: first applying regularization to the pruned components, followed by pruning and then RFT. Unlike traditional paradigms that first prune based on importance and then apply RFT, DReSS transfers critical information from the pruned parameter space to the remaining components during regularization, effectively mitigating the irreversible performance degradation caused by information loss. DReSS demonstrates superior performance in both generation and zero-shot tasks, significantly outperforming existing pruning methods. For instance, DReSS surpasses the powerful LLM Surgeon by 20% in perplexity on LLaMA2-7B. On OPT-13B, with a pruning ratio of 25%, the average accuracy drops by only 1%, while achieving a 1.41× speedup compared to the dense model. Moreover, DReSS requires only 25% of the data to achieve comparable performance to previous methods, substantially reducing both data and computational costs while minimizing the reliance on RFT.
Impact Statement
Pruning plays a crucial role in model compression. DReSS is an efficient and low-overhead pruning method, which enables smaller models to achieve competitive performance, especially in resource-constrained scenarios and it also provides some insights for the development of LLMs.
Appendix A Proofs of the Propositions
A.1 Proof of Proposition 3.1
@@ Appendix A Proofs of the Propositions
Let us first assume that $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{n\times k}$. We can express $A$ as $\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}$, where $\mathbf{a}_{i}\in\mathbb{R}^{m\times 1}$, and $B$ as $\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}$, where $\mathbf{b}_{i}\in\mathbb{R}^{k\times 1}$. Then, we obtain $C\in\mathbb{R}^{m\times k}$ as shown in Equation 11. It is important to note that the final summation in Equation 11 refers to the summation of $n$ rank-1 matrices, not a scalar summation.
$\begin{split}C=AB=\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}\times\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}=\sum_{i=1}^{n}\mathbf{a}_{i}\mathbf{b}_{i}^{T}\end{split}$ (11)
According to Equation 11, when some columns of $A$ are zero, the rows of $B$ that correspond to the outer product with the zero columns of $A$ can take any values, as the product of zero and any value is zero. These rows can therefore be set to zero.
@@
Applying the above conclusion to the parameters of a transformer layer, if certain columns of $W_{down}^{i-1}$ in the FFN layer of the $(i-1)$th transformer layer are pruned, then, since $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th transformer layer are immediately multiplied by the result of $W_{down}^{i-1}$, the corresponding rows in $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ can be set to zero. Additionally, if certain columns of $W_{o}^{i}$ in the $i$th transformer layer are pruned, the $W_{up}^{i}$ in the FFN block will immediately multiply with it, so the corresponding rows in $W_{up}^{i}$ should also be pruned. Thus, we have completed the proof of Proposition 3.1.
A.2 Proof of Proposition 3.2
@@
This transformation allows the optimization problem to be solved without directly involving the absolute value function, resulting in an equivalent constrained optimization problem that can be addressed via backpropagation. Thus, the proof of the Proposition 3.2 is complete.
Appendix B Detailed Implementation
In this part, we first introduce several hyperparameter settings, with the detailed results shown in Table 7. In our experiments, we employ FP16 precision for all evaluated models, including Phi-2, OPT-2.7B, OPT-6.7B, OPT-13B, LLaMA2-7B, and LLaMA2-13B. For all RFT configurations, we set the LoRA rank $r$ to 32, the scaling factor $\alpha$ to 10, and the sequence length to 2048. All other hyperparameters follow the default settings provided in the Hugging Face PEFT package (Mangrulkar et al., 2022). Due to limited computational resources, we set the batch size to 64. In future work, we will further explore the impact of a broader range of batch sizes on the performance of the pruning method. These choices ensure a balance between model adaptability and computational efficiency, leveraging parameter-efficient fine-tuning while maintaining stable optimization dynamics.
@@ Appendix B Detailed Implementation
Precision LoRA Rank Scaling Factor Max Sequence Length Batch Size Learning Rate Early Stop Threshold Min Delta FP16 32 10 2048 64 2e-5 5 1e-4
@@
To ensure a fair comparison between DReSS and other methods, we maintain consistency in the data used across all approaches. Specifically, the data used by DReSS for regularization, by LLM Surgeon for periodic updates of model weights and structures, by SliceGPT for selecting channel importance, and by SLEB for identifying crucial transformer layers are identical. Furthermore, we ensure that the data employed during the RFT process is consistent across all methods, thereby enabling a controlled and equitable evaluation framework. Following previous works (Ashkboos et al., 2024; Song et al., 2024), for the comparison unstructured pruning methods like Mangniude, Wanda, and SparseGPT, we ensure that the data used to compute the importance of individual weights is the same as the data used by DReSS for regularization.
Appendix C Results of Comparison with Unstructured Pruning Methods
We compare DReSS at a 25% pruning ratio with unstructured pruning methods employing 2:4 sparsity (where two out of every four consecutive parameters are set to zero) and present the experimental results in Table 8. For unstructured pruning methods, only calibration data is required. To ensure consistency, all methods utilize a total of 1,000 samples, randomly selected from WikiText-2. The models used in these experiments consists of the OPT and LLaMA model families.
Compare to structured pruning methods, the primary advantage of unstructured pruning lies in its flexibility, enabling the removal of specific weights while preserving critical information in the remaining parameters of LLMs. Based on the results in Table 8, it can be seen that DReSS with a pruning ratio of 25% significantly outperforms existing unstructured pruning methods with a sparsity of 2:4 on generative and zero-shot tasks, demonstrating the powerful performance of DReSS. The reason DReSS outperforms unstructured pruning in comparisons lies in the regularization process before pruning. This process transfers important information hidden in the parameter space to the remaining parts of the model, preventing irreversible performance degradation caused by direct pruning and the resulting loss of information.
Table 8: Performance comparison with unstructured pruning methods. ‘PPL’ refers to the perplexity on WikiText-2. The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy. The best result is highlighted in bold, and the second-best in underlined. DReSS-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while DReSS-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
Model Method Pruning Ratio PPL ($\downarrow$) PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 Magnitude 2:4 265 57.46 42.19 35.62 32.58 16.42 36.85 Wanda 2:4 16.54 64.32 57.84 41.63 43.47 23.44 46.14 SparseGPT 2:4 13.48 68.58 56.72 45.36 46.89 25.92 48.69 DReSS-$\ell_{2}$ 25% 12.52 69.12 58.98 49.10 50.62 28.15 51.20 DReSS-$\ell_{1}$ 25% 12.48 68.89 58.33 48.62 51.67 27.55 51.01 OPT-6.7B Dense 0% 10.85 76.39 65.19 67.16 60.14 34.64 60.70 Magnitude 2:4 969 62.16 48.74 43.48 41.57 20.63 43.32 Wanda 2:4 13.29 71.76 60.22 54.20 51.18 28.33 53.14 SparseGPT 2:4 11.55 74.21 60.77 57.25 53.03 29.44 54.94 DReSS-$\ell_{2}$ 25% 11.06 74.26 63.12 61.26 56.64 31.77 57.41 DReSS-$\ell_{1}$ 25% 11.74 73.86 64.28 60.45 56.36 32.07 57.40 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 Magnitude 2:4 1.2e4 64.38 52.19 49.46 44.58 21.25 46.37 Wanda 2:4 13.49 72.25 61.72 57.97 53.96 29.69 55.12 SparseGPT 2:4 11.17 74.16 62.43 59.18 56.27 31.74 56.76 DReSS-$\ell_{2}$ 25% 10.38 74.06 64.57 66.82 61.56 35.33 60.47 DReSS-$\ell_{1}$ 25% 10.56 73.65 64.38 66.46 61.15 34.74 60.07 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 Magnitude 2:4 54.59 64.42 60.93 56.72 59.18 27.31 53.71 Wanda 2:4 11.35 70.84 62.27 55.33 57.58 31.91 55.59 SparseGPT 2:4 8.67 72.14 64.96 58.93 60.90 34.22 58.23 DReSS-$\ell_{2}$ 25% 5.86 73.18 66.49 61.73 65.42 40.86 61.54 DReSS-$\ell_{1}$ 25% 5.81 73.42 65.73 61.92 65.26 39.68 61.20 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 Magnitude 2:4 8.33 70.47 62.04 50.65 62.46 31.74 55.47 Wanda 2:4 8.32 73.94 67.01 63.09 64.31 37.80 61.23 SparseGPT 2:4 7.07 75.46 68.51 65.52 66.04 39.76 63.06 DReSS-$\ell_{2}$ 25% 5.12 76.14 71.22 76.51 73.45 46.87 68.84 DReSS-$\ell_{1}$ 25% 5.16 77.24 70.61 76.75 73.84 45.46 68.78
Appendix D Changes in the Parts Without Regularization

### v3

@@ 4.6 Minimal Overhead
We evaluated the perplexity of each method on LLaMA2-7B using varying amounts of data, keeping all other conditions consistent with Section 4.2 and only change the data size from 500 to 8,000. As shown in Figure 4, when only 1,000 samples were used for regularization and RFT, DReSS outperformed the other methods that used 4,000 samples. This further highlights the effectiveness and efficiency of applying regularization prior to pruning, as it transfers critical information in advance, enabling DReSS to get strong performance with minimal data overhead.
@@ 4.7 Ablation Study
As shown in Table 5, both pruning without regularization and full parameter regularization significantly increase perplexity on WikiText-2 and reduce average accuracy at 25% sparsity. This highlights the importance of applying regularization precisely to the pruned components. Additionally, RFT after pruning has minimal impact on model performance, indicating that the RFT process is optional.
Data Ratio of Regualrization and RFT: With other settings consistent with Section 4.2, we only vary the data ratio $[4:1,3:1,2:1,1:1,1:2,1:3]$. The lowest perplexity is achieved under 3:1 ratio. We speculate that pre-pruning regularization may be more critical than post-pruning RFT.
Model Setting PPL($\downarrow$) AVG_ACC
LLaMA2-7B DReSS 5.86 61.54
w FPR 12.68 50.26
w/o R 22.38 47.25
w/o RFT 5.97 59.56
LLaMA2-13B DReSS 5.12 68.84
w FPR 10.53 53.28
w/o R 18.94 50.17
w/o RFT 5.39 67.96
Table 5: Ablation results on LLaMA2-7B, LLaMA2-13B. ‘FPR’ is full parameter regularization, ‘R’ is regularization on selected channels. ‘w/o’ means removing specific parts.
Trade-Off Between Language Modeling and Regularization: Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$. Optimal $\lambda$ is $10^{-3}$, demonstrating the importance of balancing language modeling loss and regularization loss. We listed the best $\lambda$ for each model in Appendix E.
4.8 Impact of Regularization
In Figure 5, we present the ratio of model weights and activation values after and before regularization of the selected portions. We find that in most intermediate layers, the weights and activation values after regularization are reduced to 30% of their original values, which indicates that information contained in this parameter space has decreased. We also present the changes in the unregularized part in Appendix F. The results show that this part has increased compared to before, indicating that it now contains more information showing that regularization transfers important information from the pruned parts to the remaining parts.
To evaluate regularization effectiveness, we divided into three stages: whether to regularize before pruning, whether to prune, and whether to regularize after pruning, totally 8 cases. For example,“N&P&R” denotes no regularization, pruning, then regularized. We used 25% sparsity LLaMA2-7B.“N&N&R” and “R&N&N” are the same. The 750 (0.75$\times$1,000) samples used for regularization are consistent with Section 4.2. As shown in Table 4, “R&P&N” and “N&P&R” outperform “N&P&N”, “R&P&R” is the best, showing that regularization improves performance before, after pruning, and both of them. “R&P&N” outperforms “N&P&R”, indicating that the paradigm of applying regularization before pruning may transfer important information to the remaining parts, thereby preserving model capacity.
@@ 5 Conclusion
In this paper, we propose a new pruning paradigm: apply regularization, prune, and finally RFT. Unlike previous paradigm that first prune and then apply RFT, DReSS transfers critical information from the pruned parameter space to the remaining parts during regularization, effectively mitigating the irreversible performance degradation caused by information loss. DReSS demonstrates superior performance in generation and zero-shot tasks, significantly outperforming existing methods. For instance, DReSS surpasses the powerful LLM Surgeon by 21% in perplexity on LLaMA2-7B. On OPT-13B, under 25% sparsity, the average accuracy drops by only 1%, while achieving 1.21× speedup compared to the dense model. Moreover, DReSS requires only 25% of the data to achieve comparable performance to previous methods, substantially reducing computational costs and the reliance on RFT. The new paradigm may provide insights for structured pruning in LLMs.
Appendix A Proof of the Proposition 1
@@
Let us first assume that $\mathbf{A}\in\mathbb{R}^{m\times n}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$. We can express $\mathbf{A}$ as $\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}$, where $\mathbf{a}_{i}\in\mathbb{R}^{m\times 1}$, and $\mathbf{B}$ as $\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}$, where $\mathbf{b}_{i}\in\mathbb{R}^{k\times 1}$. Then, we obtain $\mathbf{C}\in\mathbb{R}^{m\times k}$ as shown in Equation 10. It is important to note that the final summation in Equation 10 refers to the summation of $n$ rank-1 matrices, not a scalar summation.
$\begin{split}\mathbf{C}=\mathbf{AB}=\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}\times\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}=\sum_{i=1}^{n}\mathbf{a}_{i}\mathbf{b}_{i}^{T}\end{split}$ (10)
According to Equation 10, To keep the elements of $\mathbf{a}_{i}\mathbf{b}_{i}^{T}$ small, both $\mathbf{a}_{i}$ and $\mathbf{b}_{i}^{T}$ must be small. So, if regularization is applied to $\mathbf{a}_{i}$ and it should also be applied to the corresponding row $\mathbf{b}_{i}^{T}$ in $\mathbf{B}$.
@@
Applying the above conclusion to the parameters of a transformer layer, if regularization is applied to certain columns of $\mathbf{W_{down}^{i-1}}$ in the FFN layer of the $(i-1)$th transformer layer, then, since $\mathbf{W_{Q}^{i}}$, $\mathbf{W_{K}^{i}}$, and $\mathbf{W_{V}^{i}}$ in the $i$th transformer layer are immediately multiplied by the result of $\mathbf{W_{down}^{i-1}}$, regularization should also be applied to the corresponding rows in $\mathbf{W_{Q}^{i}}$, $\mathbf{W_{K}^{i}}$, and $\mathbf{W_{V}^{i}}$. Additionally, if regularization is applied to certain columns of $\mathbf{W_{o}^{i}}$ in the $i$th transformer layer, the $\mathbf{W_{up}^{i}}$ in the FFN block will immediately multiply with it, so regularization should also be applied to the corresponding rows in $\mathbf{W_{up}^{i}}$. Thus, we have completed the proof of Proposition 1.
Appendix B Proof of Proposition 2
@@
This transformation allows the optimization problem to be solved without directly involving the absolute value function, resulting in an equivalent constrained optimization problem that can be addressed via backpropagation. Thus, the proof of the Proposition 2 is complete.
Appendix C Completed Training Objective Function
The final training objective has two parts: language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$, regularization loss. Regularization loss has three parts: Attention loss, FFN loss, remain loss. In Equation 7, for clarity, we only list one FFN layer, and principles for the others are similar. By using Proposition 2, the problem can be equivalently transformed into a constrained and differentiable optimization problem, which can be directly solved using the BP algorithm. The final objective function and constraints are as follows:
$\mathcal{L}_{\text{sum}}$ $=\mathcal{L}(\mathbf{W,X})+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{att}}^{i}+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{FFN}}^{i}+\tilde{\mathcal{L}}_{\text{remain}}$ (14)
~$\tilde{\mathcal{L}}_{\text{att}}^{i}$ $=\lambda(\mathbf{\mathbf{1}^{T}Y_{3}^{i}\mathbf{1}+\mathbf{1}^{T}Y_{4}^{i}\mathbf{1}+\mathbf{1}^{T}Y_{5}^{i}\mathbf{1}+\mathbf{1}^{T}Y_{6}^{i}\mathbf{1}})$
~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ $=\lambda(\mathbf{\mathbf{1}^{T}Y_{1}^{i}\mathbf{1}+\mathbf{1}^{T}Y_{2}^{i}\mathbf{1}})$
~$\tilde{\mathcal{L}}_{\text{remain}}$ $=\lambda\Big{(}\mathbf{1}^{T}\mathbf{Y_{7}}\mathbf{1}+\mathbf{1}^{T}\mathbf{Y_{8}}\mathbf{1}+\sum_{i=1}^{l}\mathbf{1}^{T}y_{9}^{i}\mathbf{1}+\sum_{i=1}^{l}\mathbf{1}^{T}y_{10}^{i}\mathbf{1}+\mathbf{1}^{T}\mathbf{Y_{11}}\mathbf{1}\Big{)}$
[…]
$-\mathbf{Y_{4}^{i}}$ $\mathbf{\leq RW_{K}^{i}\leq Y_{4}^{i},\quad-Y_{5}^{i}\leq RW_{V}^{i}\leq Y_{5}^{i},\quad-Y_{6}^{i}\leq W_{o}^{i}R\leq Y_{6}^{i}},$
$-\mathbf{Y_{7}}$ $\mathbf{\leq W_{emb}R\leq Y_{7},\quad-Y_{8}\leq W_{pos}R\leq Y_{8}},\quad-y_{9}^{i}\mathbf{\leq R\alpha^{i}}\leq y_{9}^{i},$
$-y_{10}^{i}$ $\mathbf{\leq R\beta^{i}}\leq y_{10}^{i},\quad-\mathbf{Y_{10}\leq RW_{lm}\leq Y_{10}}$
$\mathbf{Y_{1}^{i}}$ $\mathbf{\geq 0,\quad Y_{2}^{i}\geq 0,\quad Y_{3}^{i}\geq 0,\quad Y_{4}^{i}\geq 0,\quad Y_{5}^{i}\geq 0,\quad Y_{6}^{i}\geq 0},$
$\mathbf{Y_{7}}$ $\mathbf{\geq 0,\quad Y_{8}\geq 0},\quad y_{9}^{i}\geq 0,\quad y_{10}^{i}\geq 0,\quad\mathbf{Y_{11}}\geq 0$
Appendix D Detailed Implementation
In this part, we first introduce several hyperparameter settings, with the detailed results shown in Table 6. In our experiments, we employ FP16 precision for all evaluated models, including Phi-2, OPT-2.7B, LLaMA3-8B, OPT-13B, LLaMA2-7B, and LLaMA2-13B. For all RFT configurations, we set the LoRA rank $r$ to 32, the scaling factor $\alpha$ to 10, and the sequence length to 2048. All other hyperparameters follow the default settings provided in the Hugging Face PEFT package (Mangrulkar et al. 2022). We set the batch size to 64. In future work, we will further explore a broader range of batch sizes.
LoRA Rank Scaling Factor Max Sequence Length Batch Size Learning Rate Early Stop Threshold
32 10 2048 64 2e-5 5
@@ Appendix D Detailed Implementation
To ensure a fair comparison between DReSS and other methods, we maintain consistency in the data used across all approaches. Specifically, the data used by DReSS for regularization, by LLM Surgeon for periodic updates of model weights and structures, by SliceGPT for selecting channel importance, and by SLEB for identifying crucial transformer layers are identical. Furthermore, we ensure that the data employed during the RFT process is consistent across all methods, thereby enabling a controlled and equitable evaluation framework. Following previous works (Ashkboos et al. 2024; Song et al. 2024), for the comparison unstructured pruning methods like Mangnitude, Wanda, and SparseGPT, we ensure that the data used to compute the importance of individual weights is the same as the data used by DReSS for regularization.
Appendix E Optimal $\lambda$ for Different Models
Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$.
we list $\lambda$ for the best performance of each model in the Table 7:
Model Phi-2 OPT-2.7B OPT-6.7B OPT-13B LLaMA2-7B LLaMA2-13B
$\lambda$ $10^{-3}$ $5\times 10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-3}$
Table 7: The optimal $\lambda$ for each model.
Appendix F Changes in the Parts Without Regularization